# Large Language Model Adaptation for Networking

### v1

@@ 5.4. Deep Dive
Importance of pre-trained and domain knowledge. To gain a deeper understanding of why LLM can be adapted for networking, we investigate the importance of both the pre-trained and learned domain knowledge of LLM in networking adaptation. We use Llama2-7B as the LLM for our exploration. First, we disable the pre-trained weights of Llama2 that represent its pre-trained knowledge, randomly initialize its weights and train it from scratch for each task. As depicted in Figure 13, the absence of pre-trained knowledge leads to dramatic performance decrease across all tasks. This indicates that the pre-trained knowledge indeed contains valuable common knowledge that can be shared across different tasks, making it necessary for networking adaptation.
Next, we preserve the pre-trained knowledge of Llama2 but disable the low-rank matrices that represent the learned domain-specific knowledge. As reported in Figure 13, the absence of domain knowledge also results in performance degradation on each task, which highlights the importance of NetLLM to acquire domain knowledge.
Figure 15. Comparing the performance of different LLMs adapted by NetLLM for VP and ABR, with state-of-the-art learning-based algorithms.
@@
Impacts of different sizes of LLMs. Next, we investigate the impacts of LLM sizes in the adaptation performance. We select OPT (Zhang et al., 2022a) as the foundational model for this investigation, which offers different versions with varying parameter sizes. The results are presented in Figure 16. As shown, when the parameter size exceeds 1B, the adapted OPT achieves superior or comparable performance to the advanced learning-based algorithms. However, for the ABR task, OPT-0.35B performs significantly worse than all baselines, potentially due to the limited common knowledge to generalize across tasks. This suggests that, in practice, LLMs with parameter sizes greater than 1B are suitable for networking adaptation, while those smaller than 1B may not be the optimal choices for adaptation.
@@ 6. Discussion
Multimodal LLMs. Multimodality has been actively investigated in the current landscape of LLM research (OpenAI, 2023; Team et al., 2023; Fu et al., 2023a; Li et al., 2023b). NetLLM can certainly benefit from this line of research. In fact, it already showcases its efficacy in adapting multimodal LLMs for networking, as illustrated in Figure 15. However, perhaps surprisingly, our preliminary findings indicate that multimodal LLMs do not necessarily outperform single-modal LLMs in the field of networking. This problem can be alleviated with the development of more generic LLMs (e.g., supporting more modalities).
Evaluating on other tasks. One natural question is whether NetLLM can be applied to other networking tasks beyond the three use cases presented in this paper. Actually, NetLLM’s design is independent on specific tasks, making it applicable to diverse networking scenarios. Moreover, by careful selection of three representative use cases ($\S$3), we ensure that the ideas of NetLLM can be easily applied to other networking tasks, such as bandwidth prediction (Mei et al., 2020), congestion control (Yen et al., 2023) and datacenter traffic optimization (Chen et al., 2018b). In the future, we aim to validate the effectiveness of NetLLM in adapting LLMs to address more networking problems.
@@ A.1. Details of Figure 2
The details to produce the results of “Prompt Learning” and “Token Prediction” in Figure 2 are illustrated in Figure 17 and described as follows. We use prompt learning (Liu et al., 2023d) to adapt the Llama2-7B LLM for the VP task. Specifically, we design a prompt template to encapsulate time-series viewports into textual prompts, and instruct Llama2 (Touvron et al., 2023) to generate answers for the VP task based on token prediction. We conduct our measurements on an existing immersive video viewport dataset (Jin et al., 2022). As a motivating example, we use Llama2 to predict future viewports in the next 1 second based on past viewports in the last 1 second. Following previous works (Guimard et al., 2022; Rondón et al., 2022), the viewport sampling rate is set to 5Hz, thus the future viewports and historical viewports are both 5-sample long. Note that Llama2 initially achieves poor performance without any fine-tune. Hence, we further uses OpenPrompt (Ding et al., 2022), an open-source prompt learning framework, to fine-tune Llama2 for 100000 iterations over the viewport dataset.
Note that in Figure 2 (middle) , we calculate the fraction of valid answers generated by token prediction. In our practical implementation, we consider an answer as valid if we can extract viewports from it following a series of pre-defined string parsing operations. On the other hand, an invalid answer is often the case where it contains invalid characters (e.g., unexpected punctuation) or misses some values. While designing complex rules to post-process answers can alleviate issue of invalid answers, it introduces the overhead of answer engineering (Li et al., 2023c).
@@ A.2. Details of NetLLM Implementation
We have integrated NetLLM into three existing codebases for VP (Wu et al., 2023), ABR (Xia et al., 2022), and CJS (Mao et al., 2019), and implemented the APIs in Figure 9 based on the functionalities provided in the codebases. Additional details of NetLLM implementation are explained as follows.
For the multimodal encoder, we utilize ViT (Dosovitskiy et al., 2021) to encode images, and 1D-CNN (Mao et al., 2017) to encode time-series and sequence data (e.g., historical throughputs and future chunk sizes at different bitrates in ABR). We leverage fully connected layer to extract features from scalar data (e.g., buffer occupancy in ABR), and use GNN (Wu et al., 2021; Mao et al., 2019) to process graph information (e.g., DAGs in CJS).
@@
Regarding the DD-LRNA scheme, we configure the context window for learning return distribution as $w=10$, and set the rank of low-rank matrices to be $r=128$. While additional tuning of $w,r$ may be beneficial, we empirically find that NetLLM performs well across a wide range of hyperparameter values (generally, $w\geq 10$ and $r\geq 32$ will yield good performance). Thus, we do not employ sophisticated methods to tune these hyperparameters. We also keep them fixed throughout the experiments in $\S$5.
@@ A.3. Overview of Baselines
Baselines for CJS. The following three baselines are implemented for the CJS task: Decima (Mao et al., 2019), first-in-first-out scheduling (labeled ”FIFO”) (Spark, 2024) and fair scheduling (labeled ”Fair”) (Spark, 2024). Decima (Mao et al., 2019) is a RL model for job scheduling in the distributed computing cluster, which develops a graph neural network (GNN) to efficiently process DAG information of job properties (e.g., resource demands and dependency). Both FIFO and Fair are two common scheduling algorithms used by data processing system Spark (Spark, 2024). The former schedules jobs in the order of their arrival and allocates the requested amount of resources to each job, while the latter schedules jobs in a “round robin” fashion to ensure that each job receives a roughly equal share of the cluster resources.
A.4. Details of Simulation Settings
@@
unseen setting3 450 50
unseen setting4 450 30
@@
VP simulation. As shown in Table 2, by default, we train and test each method on a large-scale viewport dataset Jin2022 (Jin et al., 2022) which consists of 2700 viewport traces recorded by 100 viewers watching 27 60-second immersive videos. The historical window (hw) and prediction window (pw) are set to be 2 seconds and 4 seconds, respectively, for the default training and testing settings. When evaluating generalization performance, we test each method on a new viewport dataset (i.e., new data distributions) and/or with a new prediction setup (i.e., increasing prediction difficulty). For instance, on unseen setting2, we evaluate each method on the new Wu2017 dataset (Wu et al., 2017). This dataset contains 9 videos with an average length of 242 seconds watched by 48 viewers, producing a total of 432 long viewport traces. As for unseen setting1, we increase pw to increase the prediction difficulty for each method. Following the setting in (Qian et al., 2018), we roughly set hw $\approx$ pw / 2 across all settings. Changing the coefficient does not qualitatively affect the results.
@@
ABR simulation. Table 3 summarizes the simulation settings for ABR. By default, we train and test all methods to stream the Envivio-Dash3 video from the DASH-246 JavaScript reference client (Form, 2016), whose format follows the GENET (Xia et al., 2022) and Pensieve (Mao et al., 2017) setting. We use the broadband FCC (Commission, 2016) traces as the default bandwidth dataset, which contains more than 90 hours of bandwidth traces. To simulate environments for generalization testing, we follow the method in Pensieve (Mao et al., 2017) to generate a synthetic video SynthVideo which shares a similar format of Envivio-Dash3 but with a larger video bitrate. Besides, we also generate a new bandwidth dataset SynthTrace according to the method in Pensieve (Mao et al., 2017), which exhibits a larger bandwidth range and more dynamic fluctuation patterns than FCC.
@@
CJS simulation. Table 4 provides the detailed information of the CJS simulation. Following Decima (Mao et al., 2019), we simulate different workload traces using a real-world dataset TPC-H (Chiba and Onodera, 2016) which contains job requests of large data volumes, high executor demands, and high degree of complexity. According to the repository of Decima (Mao et al., 2019), we set the number of job requests to be 200 and the number of executor resources (representing computation resources) to be 50k units as the default training and testing settings. To evaluate the generalization performance of each method, we simulate various unseen harder workloads by increasing the number of job requests and reducing the number of executor resources, as also done in Decima (Mao et al., 2019).
Note. It is worth noting that while the default training and testing environments share the same setting, no testing samples are directly exposed in the training environment. For instance, in ABR task, we split the bandwidth dataset into training and testing sets so that the testing bandwidth traces are different from the training ones.
@@ A.5. Real-world ABR Testbed Setup
We leverage the testbed from GENET (Xia et al., 2022) to test the NetLLM-adapted Llama2 in real-world client-server ABR system. The testbed modifies dash.js (version 2.4) to support BBA, MPC and GENET streaming algorithms. We further modify the dash.js to support the adapted Llama2. In our real-world tests, the client video player is a Google Chrome browser (version 87) and the video server (Apache version 2.7) runs on the same machine as the client. All tests are performed on our Linux server, with two different ports to emulate the ABR client and video server. We then use Mahimahi (Netravali et al., 2015) to emulate different network environments from the broadband traces (Commission, 2016) and cellular mobile traces (Riiser et al., 2013), along with an 80ms RTT, between the client and server.

### v3

@@ 5.4. Deep Dive
Importance of pre-trained and domain knowledge. To gain a deeper understanding of why LLMs can be adapted for networking, we investigate the importance of both the pre-trained and learned domain knowledge of LLMs in networking adaptation. We use Llama2-7B as the LLM for our exploration. First, we disable the pre-trained weights of Llama2 that represent its pre-trained knowledge, randomly initialize its weights and train it from scratch for each task. As depicted in Figure 13, the absence of pre-trained knowledge leads to dramatic performance drop across all tasks. This indicates that while LLMs are pre-trained over text corpora to acquire language knowledge, their emergent abilities (e.g., planning (Xi et al., 2023), pattern mining (Zitkovich et al., 2023)) are indeed universal and applicable across domains, including networking. For instance, the pattern mining ability of LLM can be utilized to mine complex changing patterns of viewport movement for accurate viewport prediction. Hence, the pre-trained knowledge of LLMs is crucial for networking adaptation.
Next, we preserve the pre-trained knowledge of Llama2 but disable the low-rank matrices that represent the learned domain knowledge. As reported in Figure 13, the absence of domain knowledge also results in performance degradation on each task, which highlights the importance of NetLLM to acquire domain knowledge.
@@
Figure 15. Comparing the performance of different LLMs adapted by NetLLM for VP and ABR with learning-based algorithms.
Impacts of different sizes of LLMs. Next, we investigate the impacts of LLM sizes on the adaptation performance. We select OPT (Zhang et al., 2022a) as the foundational model for this investigation, which offers different versions with varying parameter sizes. The results are presented in Figure 16. As shown, when the parameter size exceeds 1B, the adapted OPT achieves superior or comparable performance to the advanced learning-based algorithms. However, for the ABR task, OPT-0.35B performs significantly worse than all baselines, potentially due to the limited common knowledge to generalize across tasks. This suggests that, in practice, LLMs with parameter sizes greater than 1B are suitable for networking adaptation, while those smaller than 1B may not be the optimal choices for adaptation.
@@ 6. Discussion
Q1: What considerations are needed when adapting LLMs for specific networking tasks using NetLLM?
@@
Q2: How does NetLLM compare to retrieval-augmented generation (RAG)?
Retrieval-augmented generation (RAG) (Pan et al., 2024; Lewis et al., 2020; Wu et al., 2022) has been recently proposed to enhance the capabilities of LLMs. RAG constructs a large corpus as an external knowledge base to store domain-specific knowledge. During the inference stage, relevant information is retrieved from the knowledge base and attended to the input context of LLMs to provide domain knowledge. Although RAG has shown efficacy in improving the performance of LLMs in NLP tasks, it faces challenges when applied in the networking field. This is because, unlike NLP where knowledge can be stored explicitly in the textual form (e.g., the names of presidents of different countries), representing domain knowledge in networking as plain texts is challenging due to the abstract and implicit nature of networking knowledge (e.g., the ABR policy to dynamically adjust bitrates based on the changing network conditions). As a result, constructing an external knowledge base for RAG is rather challenging in the field of networking. In contrast, NetLLM takes a different approach by designing an efficient DD-LRNA module, which enables LLMs to automatically and effectively learn domain-specific knowledge for networking. With this module, NetLLM facilitates efficient acquisition of networking knowledge without relying on the construction of an external knowledge base.
Q3: How to reduce the computation overhead of LLMs?
[…]
Q4: Why can LLMs be useful in networking?
In $\S$5.4, we have identified from the high-level perspective that the pre-trained knowledge of LLMs is one of the dominant factors to their success in networking. However, further investigations into the internal working mechanisms of LLMs are crucial to improve their explainability. Gaining deeper insights into the explainability of LLMs enables researchers to comprehend their capabilities, limitations, and areas for improvement (Zhao et al., 2024; Alammar, 2021). This understanding, in turn, paves the way for the development of more reliable and secure LLM-based networking systems that can be trustfully deployed in real-world scenarios. Therefore, a significant future research direction lies in designing an interpretable system to elucidate the behaviors of LLMs in the context of networking, which will greatly facilitate the effective utilization of LLMs in networking.
@@ 7. Concluding Remarks
ACKNOWLEDGEMENTS
We thank the anonymous SIGCOMM reviewers and our shepherd, Shay Vargaftik, for their invaluable feedbacks. This work was supported in part by NSFC with Grant No. 62293482, the Basic Research Project No. HZQB-KCZYZ-2021067 of Hetao Shenzhen-HK S&T Cooperation Zone, NSFC with Grant No. 62102342, the Shenzhen Science and Technology Program with Grant No. RCBS20221008093120 047, the Young Elite Scientists Sponsorship Program of CAST with Grant No. 2022QNRC001, the Shenzhen Outstanding Talents Training Fund 202002, the Guangdong Research Projects No. 2017ZT07X1 52 and No. 2019CX01X104, the Guangdong Provincial Key Laboratory of Future Networks of Intelligence (Grant No. 2022B1212010001), and the Shenzhen Key Laboratory of Big Data and Artificial Intelligence (Grant No. ZDSYS201707251409055). Zhi Wang’s work was supported in part by Shenzhen Science and Technology Program (Grant No. JCYJ20220818101014030). Junchen Jiang’s work was supported in part by NSF CNS 2146496, 1901466, 2313190, 2131826.
@@ A.1. Details of Figure 2
The details to produce the results of “Prompt Learning” and “Token Prediction” in Figure 2 are illustrated in Figure 17 and described as follows. We use prompt learning (Liu et al., 2023d) to adapt the Llama2-7B LLM for the VP task. Specifically, we design a prompt template to encapsulate time-series viewports into textual prompts, and instruct Llama2 (Touvron et al., 2023) to generate answers for the VP task based on token prediction. We conduct our measurements on an existing immersive video viewport dataset (Jin et al., 2022). As a motivating example, we use Llama2 to predict future viewports in the next 1 second based on past viewports in the last 1 second. Following previous works (Guimard et al., 2022; Rondón et al., 2022), the viewport sampling rate is set to 5Hz, thus the future viewports and historical viewports are both 5-sample long. Note that Llama2 initially achieves poor performance without any fine-tuning. Hence, we further use OpenPrompt (Ding et al., 2022), an open-source prompt learning framework, to fine-tune Llama2 for 100000 iterations over the viewport dataset.
Note that in Figure 2 (middle), we calculate the fraction of valid answers generated by token prediction. In our practical implementation, we consider an answer as valid if we can extract viewports from it following a series of pre-defined string parsing operations. On the other hand, an invalid answer is often the case where it contains invalid characters (e.g., unexpected punctuation) or misses some values. While designing complex rules to post-process answers can alleviate the issue of invalid answers, it introduces the overhead of answer engineering (Li et al., 2023d).
@@ A.2. Details of NetLLM Implementation
We have integrated NetLLM into three existing codebases for VP (Wu et al., 2023), ABR (Xia et al., 2022), and CJS (Gertsman, 2024). We have implemented the APIs in Figure 9 based on the functionalities provided in the codebases. Additional details of NetLLM implementation are explained as follows.
For the multimodal encoder, we utilize ViT (Dosovitskiy et al., 2021) to encode images, and 1D-CNN (Mao et al., 2017) to encode time-series and sequence data (e.g., historical throughputs and future chunk sizes at different bitrates in ABR). We leverage fully connected layer to extract features from scalar data (e.g., buffer occupancy in ABR), and use GNN (Wu et al., 2021; Mao et al., 2019) to process graph information (e.g., DAGs in CJS). By default, the multimodal encoders are trainable, except that the parameters of ViT are frozen. This is because ViT has open source pre-trained weights which can be used to effectively extract image features.
@@
Regarding the DD-LRNA scheme, we configure the context window $w$ for learning return distribution as 10 and 20 for ABR and CJS, respectively. We then set the rank $r$ of low-rank matrices to be 32, 128 and 128, for VP, ABR and CJS, respectively. While additional tuning of $w,r$ may be beneficial, we empirically find that NetLLM performs well across a wide range of hyperparameter values (generally, $w\geq 10$ and $r\geq 32$ will yield good performance). Thus, we do not employ sophisticated methods to tune these hyperparameters and keep them fixed throughout the experiments in $\S$5. As for experience collection, we use GENET (Xia et al., 2022) and Decima (Mao et al., 2019) to collect experience datasets for the RL-based ABR and CJS tasks, respectively. While using more algorithms to interact with the environments for more epochs to expend the datasets may yield potential benefits, we leave this for future exploration.
@@ A.3. Overview of Baselines
Note that since the open-source codes of TRACK are originally written in Keras (Gulli and Pal, 2017), we carefully convert its codes into PyTorch (Paszke et al., 2019) to make TRACK compatible with the VP codebase (Wu et al., 2023). We have ensured that our implementation preserves the same functionality of TRACK as its original implementation. Besides, as TRACK does not offer pre-trained model weights, we re-train it from scratch on our datasets with the same training hyperparameters described in TRACK’s paper and codes. As for rule-based algorithms LR and Velocity, which do not provide open source implementation, we implement them ourselves by strictly following the same ideas and formulas presented in their respective papers.
To implement the aforementioned ABR baselines, we utilize the open-source codes of GENET (Xia et al., 2022), which already include the implementation of the three algorithms. Furthermore, we re-use the pre-trained model weights of GENET [^5] https://github.com/GenetProject/Genet/tree/main/src/emulator/abr/pensieve/data/mahimahi_new_best_models/ADR_model for our experiments.
@@
Baselines for CJS. The following three baselines are implemented for the CJS task: Decima (Mao et al., 2019), first-in-first-out scheduling (labeled ”FIFO”) (Spark, 2024) and fair scheduling (labeled ”Fair”) (Spark, 2024). Decima (Mao et al., 2019) is a RL model for job scheduling in the distributed computing cluster, which develops a graph neural network (GNN) to efficiently process DAG information of job properties (e.g., resource demands and dependency). Both FIFO and Fair are two common scheduling algorithms used by data processing system Spark (Spark, 2024). The former schedules jobs in the order of their arrival and allocates the requested amount of resources to each job, while the latter schedules jobs in a “round robin” fashion to ensure that each job receives a roughly equal share of the cluster.
We utilize the PyTorch re-implementation of Decima (Gertsman, 2024) for our experiments as the original implementation is somewhat outdated. Furthermore, we make use of the pre-trained model weights of Decima [^6] https://github.com/ArchieGertsman/spark-sched-sim/tree/main/models/decima provided in (Gertsman, 2024). Additionally, we adopt the implementation of FIFO and Fair from the same source (Gertsman, 2024).
A.4. Details of Simulation Settings
@@
unseen setting2 450 50
unseen setting3 450 30
@@
VP simulation. As shown in Table 2, by default, we train and test each method on a large-scale viewport dataset Jin2022 (Jin et al., 2022) which records the viewport traces from 84 viewers [^7] The Jin2022 dataset originally contains the viewport traces from 100 viewers (Jin et al., 2022). We filter out those incomplete ones (i.e., less than 60 seconds in duration) and finally use the traces from 84 viewers for experiments. watching 27 60-second immersive videos. We randomly select 15 videos and 42 viewers for training, 6 videos and 21 viewers for validation, 6 videos and 21 viewers for testing. This results in a total of 882 traces for experiments. The historical window (hw) and prediction window (pw) are set to be 2 seconds and 4 seconds, respectively, for the default training and testing settings.
@@
When evaluating generalization performance, we test each method on a new viewport dataset (i.e., new data distributions) and/or with a new prediction setup (i.e., increasing prediction difficulty). For instance, on unseen setting2, we evaluate each method on the new Wu2017 dataset (Wu et al., 2017). This dataset contains 9 videos [^8] The Wu2017 dataset originally includes 18 videos. We use the first 9 videos as viewers are free to look around when watching these videos. with an average length of 242 seconds watched by 48 viewers. We randomly sample 4 videos and 9 viewers from the dataset, resulting in 36 long viewport traces for testing generalization. As for unseen setting1, we increase pw to increase the prediction difficulty for each method. Following the setting in (Qian et al., 2018), we roughly set hw $\approx$ pw / 2 across all settings. Changing the coefficient does not qualitatively affect the results.
@@
ABR simulation. Table 3 summarizes the simulation settings for ABR. By default, we train and test all methods to stream the Envivio-Dash3 video from the DASH-246 JavaScript reference client (Form, 2016), whose format follows the GENET (Xia et al., 2022) and Pensieve (Mao et al., 2017) setting. We use the broadband FCC (Commission, 2016) traces as the default bandwidth dataset. In particular, we use the same traces for training and validation as those used by GENET, which compromise 235 traces for training and 150 traces for validation. Then, we randomly sample 100 traces from the remaining dataset for testing. This results in the use of more than 90 hours of bandwidth traces for experiments. To simulate environments for generalization testing, we follow the method in Pensieve (Mao et al., 2017) to generate a synthetic video SynthVideo which shares a similar format of Envivio-Dash3 but with a larger video bitrate. Besides, we also generate a new bandwidth dataset SynthTrace with 100 traces according to the method in Pensieve (Mao et al., 2017), which exhibits a larger bandwidth range and more dynamic fluctuation patterns than FCC.
@@
CJS simulation. Table 4 provides the detailed information of the CJS simulation. Following Decima (Mao et al., 2019), we simulate different workload traces using a real-world dataset TPC-H (Chiba and Onodera, 2016) which contains job requests of large data volumes, high executor demands, and high degree of complexity. To be consistent with the settings used by the pre-trained Decima in (Gertsman, 2024), we set the number of job requests to be 200 and the number of executor resources (representing computation resources) to be 50k units as the default training and testing settings. To evaluate the generalization performance of each method, we simulate various unseen harder workloads by increasing the number of job requests and reducing the number of executor resources, as also done in Decima (Mao et al., 2019). Note that in each setting the job requests are randomly sampled from the TPC-H dataset. Besides, when evaluating on the default testing setting, we have ensured that the job requests are different from those in the training setting. This can be easily done by setting different random seeds for data sampling.
@@ A.5. Real-world ABR Testbed Setup
We leverage the testbed from GENET (Xia et al., 2022) to test the NetLLM-adapted Llama2 in a real-world client-server ABR system. The testbed modifies dash.js (version 2.4) to support BBA, MPC and GENET streaming algorithms. We further modify the dash.js to support the adapted Llama2. In our real-world tests, the client video player is a Google Chrome browser (version 87) and the video server (Apache version 2.7) runs on the same machine as the client. All tests are performed on our Linux server, with two different ports to emulate the ABR client and video server. We then use Mahimahi (Netravali et al., 2015) to emulate different network environments from the broadband traces (Commission, 2016) and cellular mobile traces (Riiser et al., 2013), along with an 80ms RTT, between the client and server. In particular, we randomly sample 100 traces from both the broadband and cellular mobile bandwidth datasets for network environment emulation.
A.6. Evaluation Metrics
Metric for VP. We use mean absolute error (MAE) as the evaluation metric for the VP task. Let $\boldsymbol{v}_{p}=(\alpha_{t},\beta_{t},\zeta_{t})$ denote the viewport coordinate at timestep $t$, where $\alpha,\beta,\zeta$ represent the roll, pitch and yaw values, respectively. Given $\boldsymbol{v}^{p}_{t},\boldsymbol{v}^{g}_{t}$ as the predicted and ground-truth viewports and $H$ as the prediction horizon, the MAE is calculated by:
$MAE=\frac{1}{H}\sum_{t=1}^{H}\frac{|\alpha_{t}^{p}-\alpha_{t}^{g}|+|\beta_{t}^{p}-\beta_{t}^{g}|+|\zeta_{t}^{p}-\zeta_{t}^{g})|}{3}$
Metric for ABR. We use quality of experience (QoE) as the evaluation metric for the ABR task, which is defined as the weighted linear combination of three metrics (Mao et al., 2017; Xia et al., 2022):
$QoE=\frac{\sum_{i=1}^{C}(Bitrate_{i}-\lambda Rebuf_{i}-\gamma BitrateChange_{i})}{C}$
where $Bitrate_{i}$ is the bitrate in Mbps of chunk $i$, $Rebuf_{i}$ is the rebuffering time in seconds of downloading chunk $i$, $BitrateChange_{i}$ is the bitrate change in Mbps between consequtive chunks, $C$ is the number of chunks of the video and $\lambda,\gamma$ are the weight parameters. Following Pensieve (Mao et al., 2017), we set $\lambda=4.3$ and $\gamma=1$.
Metric for CJS. We use job completion time (JCT) (Mao et al., 2019) as the evaluation metric for the CJS task. Let $t_{s}$ denote the arrival time of a job and $t_{e}$ denote the finishing time of a job. The JCT is calculated by:
$JCT=t_{e}-t_{s}$